# Recurrent neural network analysis of single trajectories switching between anomalous diffusion states

### v1

@@ C.1 Confidence levels for the CPDA
Figure 11: Comparison of CPDA and $\mathtt{sequandi}$ Performance. The confidence levels used are from  C.1. (a) shows the RMSE of different confidence levels for CPDA and Sequandi, while (b) presents the JSC performance under the same settings. Lower RMSE and higher JSC values indicate better performance.
@@
Appendix D Benchmark dataset experiment-wise performance
@@
In this appendix we report the performance of the segment-wise predictions on the benchmark dataset discussed in Section 5.4 for each experiment.
@@
Figure 13: Ground truth vs. predicted anomalous diffusion exponent. For each experiment in the benchmark dataset, a heat-map of the predicted segment-wise $\hat{\alpha}_{i}$ and their respective ground truths.
@@
Figure 14: Ground truth vs. predicted generalised diffusion coefficient. For each experiment in the benchmark dataset, a heat-map of the predicted segment-wise $\hat{K}_{i}$ and their respective ground truths. The x and y axis scales are logarithmic, and the colourisation of the heat-map is with a logarithmic normalisation due to the more dilute grid.

### v2

@@ D.1 Confidence levels for the CPDA
Figure 13: Comparison of CPDA and $\mathtt{sequandi}$ Performance. The confidence levels used are from  D.1. (a) shows the RMSE of different confidence levels for CPDA and Sequandi, while (b) presents the JSC performance under the same settings. Lower RMSE and higher JSC values indicate better performance.
@@
Appendix E Visualising sequandi performances on example trajectories
E.1 Comparison with TA-MSD and Ref. [55]
Figure 15: Example of $\hat{\alpha}(t)$ and $\hat{K}(t)$ and ground truth for the trajectory of Fig. 1 repeated twice. (a) The orange line is sequandi’s point-wise prediction of the anomalous exponent $\hat{\alpha}(t)$, and the blue piece-wise constant line is the ground truth $\alpha(t)$. The grey line is the prediction with the method from Ref. [55], and the brown line is the TA-MSD prediction, computed with a sliding window of length 50. (b) The orange line is based on sequandi’s point-wise prediction of the generalised diffusion coefficient $\hat{K}(t)$, and is $\log(\hat{K}(t)+1)$. The blue piece-wise constant line is the ground truth $\log(K(t)+1)$, and the brown line is the TA-MSD prediction, computed with a sliding window of length 50.
A standard way to characterise anomalous diffusion from single trajectories is to compute the TA-MSD [42, 43, 44]. If the process considered is ergodic [24], which is the case for the fractional Brownian motion (fBm) trajectories considered here, plotting the TA-MSD in a log-log plot yields approximately a straight line where the slope is the anomalous diffusion exponent, and the offset is the generalised diffusion coefficient. We use the first 5 lag times to fit the straight line, as in Ref. [55]. A more refined method, with excellent performance on fBm switching between different exponents (one switch), was derived in Ref. [55]. In Fig. 3, we compare sequandi to these traditional methods for the challenging case of multiple switches. Since the method of Ref. [55] requires a segment of at least 256 points, we apply it to two consecutive copies of the trajectory of Fig. 1, which gives a trajectory containing 400 data points, switching between different diffusive behaviours as shown in Fig. 15. In Fig. 3 in the main text, we report the results on the first 200 points of the trajectory. In Fig. 15 in this Appendix, we report the performance on the duplicated trajectory, containing 400 points. Even though sequandi was trained on trajectories of length 200, it can easily generalise to longer trajectories and greatly outperform the traditional methods. For the original first 200 points, sequandi tracks the ground truth with an $\textrm{MAE}_{\alpha(t)}=0.134$ and $\textrm{MSLE}_{K(t)}=0.016$. Using the repeated trajectory of 400 frames, sequandi achieves a comparable and even better performance, $\textrm{MAE}_{\alpha(t)}=0.116$ and $\textrm{MSLE}_{K(t)}=0.011$. The RNN from Ref. [55] has $\textrm{MAE}_{\alpha(t)}=0.248$, and the TA-MSD curve has $\textrm{MAE}_{\alpha(t)}=0.280$ and $\textrm{MSLE}_{K(t)}=0.386$.
E.2 Additional example trajectories
To complement the main text, in this appendix we provide more examples of point-wise and CP location inference on individual trajectories, generated from the same ensemble as the test dataset in Fig. 11. The trajectories are longer than 256 frames, to accommodate the use of the Ref. [55] network without needing to repeat them. The $x$ and $y$ trajectories of the two new trajectories, (a) and (b) in Fig. 16, are plotted in the top row. (a) features an intermediate and long duration segment of different state parameters, with short switches in between. (b) has rapid hopping throughout, visiting a trapped, immobile state.
Figure 16: Inference of diffusion parameters and segments on two additional trajectories (a) and (b) Top row: $x$ and $y$ coordinate tracks of the diffusive trajectories, with the ground truth changepoint times $t^{\textrm{CP}}$ denoted by the dashed vertical line. Middle row: anomalous diffusion exponent evolution with ground truth $\alpha(t)$ in blue, and point-wise predictions $\hat{\alpha}(t)$ in orange. Predicted changepoint times $\hat{t}^{\textrm{CP}}$ given by the solid black vertical lines. The grey line is the prediction with the method from Ref. [55], and the brown line is the TA-MSD prediction, computed with a sliding window of length 50. Bottom row: likewise for the generalised diffusion coefficient $K(t)$. For the sequandi predictions in (a), $\textrm{MAE}_{\alpha(t)}=0.132$ and $\textrm{MSLE}_{K(t)}=0.040$. For (b), $\textrm{MAE}_{\alpha(t)}=0.079$ and $\textrm{MSLE}_{K(t)}=0.045$.
sequandi manages to track large jumps in $K$, with inference of both parameters improving with longer segment duration. However, we highlight once more that these are very frequent switching of diffusion states, and trajectories overall have a short duration. sequandi detects accurately the changepoints in both trajectories and wrongly identifies a false positive in trajectory (b). Comparing sequandi with the network of [55], which struggles on these short segments and the TA-MSD computed with sliding windows of total length 50, which coarsely tracks the varying diffusion parameters confirms the great improvement provided by sequandi for trajectories featuring frequent switches.
Appendix F Benchmark dataset experiment-wise performance
@@
In this Appendix, we report the performance of the segment-wise predictions on the benchmark dataset discussed in Section 5.4 for each experiment.
@@
Figure 17: Ground truth vs. predicted anomalous diffusion exponent. For each experiment in the benchmark dataset, a density map of the predicted segment-wise $\hat{\alpha}_{i}$ and their respective ground truths.
@@
Figure 18: Ground truth vs. predicted generalised diffusion coefficient. For each experiment in the benchmark dataset, a density map of the predicted segment-wise $\hat{K}_{i}$ and their respective ground truths. The x and y axis scales are logarithmic, and the colourisation of the density map is with a logarithmic normalisation due to the more dilute grid.